# Primal-Dual iLQR

### 01 Mar 2024 v1

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual SQP problem. As our algorithm is a specialization of NPSQP (Gill et al. 1992), it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS (Giftthaler et al. 2018) and FDDP (Mastalli et al. 2020).
@@ II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP (Mayne 1966), achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in (Shoemaker and Liao 1991). iLQR (Todorov and Li 2004) can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton (Dunn and Bertsekas 1989) exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
@@
DDP, iLQR, and Stagewise Newton are all single shooting methods. This makes them hard to warm start, as the state trajectory cannot be chosen independently of the control trajectory. This can result in more iterations being required for these methods to converge, as well as higher likelihoods of convergence to undesired local optima.
@@
(Giftthaler et al. 2018) introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. (Mastalli et al. 2020) introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
@@
Much work has also been done in handling discrete-time optimal control problems with stagewise constraints. ALTRO (Howell, Jackson, and Manchester 2019), FATROP (Vanroye et al. 2023), and acados (Verschueren et al. 2022) are likely the most competitive open source software packages for this application at the time of writing. ALTRO handles constraints via the augmented Lagrangian method, while FATROP uses an interior point method and acados uses an SQP approach. ForcesNLP (Zanelli et al.) is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
(Singh, Slotine, and Sindhwani 2022) introduces a family of methods combining SQP and DDP, one of which using a similar approach to the one discussed in this paper (what the authors call the “open loop (OL)” variant), which seems to perform poorly in their benchmarks. At several points in their paper and code (available in the Trajax repository on GitHub), there are missing affine terms, for example in the linearized dynamics, which would likely explain the benchmark results.
@@
At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT (Wächter and Biegler 2006) with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$. Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in (Rao, Wright, and Rawlings 1998). The modified KKT system will only have the extra stagewise affine constraints when the problem being solved has other stagewise equality constraints besides the dynamics. A similar algorithm that does not handle other general equality constraints had previously been presented in (Pavlov, Shames, and Manzie 2020). As compared to the specialization of FATROP to problems without stagewise equality constraints, our method has the advantage of not requiring second order corrections for fast local convergence. ForcesNLP also follows an interior point approach, but relies on a Cholesky factorizations instead of LQR decompositions for solving its linear systems.
@@
On the other hand, the augmented Lagrangian technique employed by ALTRO yields linear systems that can be solved via the LQR algorithm, even in the presence of stagewise equality constraints. However, ALTRO has a slightly unorthodox way of doing multiple shooting: it modifies the dynamics $x_{n+1}=f_{n}(x_{n},u_{n})$ into $x_{n+1}=f_{n}(x_{n},u_{n})+e_{n}$ and adds the constraints $e_{n}=0$. This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting method. This approach has the disadvantage of substantially expanding the control dimension.
@@
Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM (Frison and Diehl 2020), an efficient interior point QP solver that exploits the discrete-time optimal control problem structure. While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to (Vanroye et al. 2023). Moreover, due to not employing a globalization strategy, global convergence is not guaranteed. When the functions defining the problem are expensive to evaluate, SQP methods become time-competitive, as the higher number of linear system solves will be compensated by the lower number of evaluations of the nonlinear problem.
@@ II-C Contributions
This framework can be extended to also handle constrained problems, using augmented Lagrangian (Howell, Jackson, and Manchester 2019), interior point (Vanroye et al. 2023), or SQP (Verschueren et al. 2022) methods. ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses). Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
@@ III-C Merit Functions
Merit functions offer a measure of progress towards the final solution and are typically used in combination with a line search procedure. Note that it would be inadequate to simply apply line search on the original objective, as SQP may (and often does) provide a search direction that locally increases that objective (usually trading that off for a reduction in the constraint violations). Filter methods, used in (Wächter and Biegler 2006), provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
@@
Note that (Gill et al. 1992) suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
@@ III-D Line Search
Note that NPSQP (Gill et al. 1992) uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
@@ VI Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step.
@@ VII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm. This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in (Gill et al. 1992).

### 05 Mar 2024 v2

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual SQP problem. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS [3] and FDDP [6].
@@ II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP [7], achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in (Shoemaker and Liao 1991). iLQR [11] can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton [1] exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
@@
DDP, iLQR, and Stagewise Newton are all single shooting methods. This makes them hard to warm start, as the state trajectory cannot be chosen independently of the control trajectory. This can result in more iterations being required for these methods to converge, as well as higher likelihoods of convergence to undesired local optima. Moreover, they require that the dynamics be evaluated sequentially, making them harder to parallelize.
@@
[3] introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. [6] introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
@@
Much work has also been done in handling discrete-time optimal control problems with stagewise constraints. ALTRO [5], FATROP [12], and acados [13] are likely the most competitive open source software packages for this application at the time of writing. ALTRO handles constraints via the augmented Lagrangian method, while FATROP uses an interior point method and acados uses an SQP approach. ForcesNLP [15] is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
@@
At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT [14] with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$. Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in [9]. The modified KKT system will only have the extra stagewise affine constraints when the problem being solved has other stagewise equality constraints besides the dynamics. A similar algorithm that does not handle other general equality constraints had previously been presented in [8]. As compared to the specialization of FATROP to problems without stagewise equality constraints, our method has the advantage of not requiring second order corrections for fast local convergence. ForcesNLP also follows an interior point approach, but relies on a Cholesky factorizations instead of LQR decompositions for solving its linear systems.
@@
On the other hand, the augmented Lagrangian technique employed by ALTRO yields linear systems that can be solved via the LQR algorithm, even in the presence of stagewise equality constraints. However, ALTRO has a slightly unorthodox way of doing multiple shooting: it modifies the dynamics $x_{n+1}=f_{n}(x_{n},u_{n})$ into $x_{n+1}=f_{n}(x_{n},u_{n})+e_{n}$ and adds the constraints $e_{n}=0$. This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting algorithm. This approach has the disadvantage of substantially expanding the control dimension.
@@
Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM [2], an efficient interior point QP solver that exploits the discrete-time optimal control problem structure. While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to [12]. Moreover, due to not employing a globalization strategy, global convergence is not guaranteed. When the functions defining the problem are expensive to evaluate, SQP methods become time-competitive, as the higher number of linear system solves will be compensated by the lower number of evaluations of the nonlinear problem.
@@ II-C Contributions
This framework can be extended to also handle constrained problems, using augmented Lagrangian [5], interior point [12], or SQP [13] methods. ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses). Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
@@ III-C Merit Functions
Merit functions offer a measure of progress towards the final solution and are typically used in combination with a line search procedure. Note that it would be inadequate to simply apply line search on the original objective, as SQP may (and often does) provide a search direction that locally increases that objective (usually trading that off for a reduction in the constraint violations). Filter methods, used in  [14], provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
@@
Note that [4] suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
@@ III-D Line Search
Note that NPSQP [4] uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
@@ VI Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. This seems to be the problem used in [16]. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step. A JAX implementation of our algorithm takes 46ms to finish solving this problem. Below, you can find a visual description of the problem and solution, as well as solver logs.
Figure 1: The solution of the quad-pendulum problem.
@@ VII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm. This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in [4].
Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton. Moreover, it allows that the evaluation of the dynamics and derivatives be computed in parallel.